1. Exhibit 22-13 Quantity of Output (Q) Total Fixed Cost (TFC) Average Fixed Cost (AFC) Total...

1.

Exhibit 22-13

Quantity of Output

(Q)

Total Fixed Cost (TFC)

Average Fixed Cost (AFC)

Total Variable Cost
(TVC)

Average Variable Cost
(AVC)

Total Cost
(TC)

Average Total Cost
(ATC)

Marginal Cost
(MC)

0

$200

$0

$200

1

$200

(A)

30

(H)

230

(M)

(S)

2

$200

(B)

50

(I)

250

(N)

(T)

3

$200

(C)

(F)

$26.67

(K)

(P)

(U)

4

$200

(D)

130

(J)

330

(Q)

(V)

5

$200

(E)

(G)

$40

(L)

(R)

(W)

Refer to Exhibit 22-13.  What dollar amounts go in blanks (G) and (H), respectively?

a.

$100; $30

b.

$400; $50

c.

$200; $30

d.

$40; $20

e.

There is not enough information to answer this question.
Answer #1

Answer
G is a total variable cost of 5 units of output
total variable cost =AVC*Q
=40*5=$200
======
H is the average variable cost of 1 unit
average variable cost=total variable cost/quantity
=30/1
=$30

option c